OpenStudy (anonymous):

Use rational exponents to write ∛x * ∜2x as a single radical expression.

OpenStudy (anonymous):

I got this far\[x ^{1/3}*2^{1/4}*x ^{1/4}\] (I'm not sure if I'm doing this right :(

OpenStudy (mertsj):

That's very good. When you are multiplying, if the bases are the same, you add the exponents.

OpenStudy (anonymous):

\[\sqrt[12]{8x^7}\]

OpenStudy (anonymous):

12rad(x)^7 4rad(2)

OpenStudy (mertsj):

\[x ^{1/3}x ^{1/4}=x ^{1/3+1/4}=x ^{7/12}\]

OpenStudy (mertsj):

also 2^1/4 =2^3/12 so we have a 12th root

OpenStudy (mertsj):